AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 302     Entry time: Sun Feb 27 11:00:08 2022
Author: RSS TD 
Subject: Current status 

12:00: 95K-->120K

12:07: 120K --> 95K

Ring current ~6e7 particles/s

Merger, Tape server (WS3): 1414 kB/s writing ok

Leakage current (WS4): 3.43 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 12-xy-2022-02-27_12-16-19.png  271 kB  | Hide | Hide all
12-xy-2022-02-27_12-16-19.png
Attachment 2: 12-yrate-2022-02-27_12-15-42.png  157 kB  | Hide | Hide all
12-yrate-2022-02-27_12-15-42.png
Attachment 3: 12-xrate-2022-02-27_12-15-03.png  156 kB  | Hide | Hide all
12-xrate-2022-02-27_12-15-03.png
ELOG V3.1.4-unknown